# Parabolic induction in the homotopy category of pro-$p$ Iwahori-Hecke modules

### v2

@@ 2.4. Recollections on abelian model categories
By combining 2.14 and 2.15, we see that for ${\mathfrak{m}},\mathfrak{n}\in{\operatorname{Mod}}(R)$ with ${\mathfrak{m}}$ Gorenstein projective and for $\theta:\mathfrak{p}\to\mathfrak{n}$ as above, $[{\mathfrak{m}},\mathfrak{n}]_{R}\cong{\operatorname{coker}}(\theta\circ(-))\cong{\operatorname{Ext}}^{1}_{S}({\mathfrak{m}},\ker(\theta))$. This generalises as follows. We say that an abelian model structure $({\mathcal{C}},{\mathcal{F}},{\mathcal{W}})$ on an abelian category ${\mathcal{A}}$ is projective, resp. injective, if ${\mathcal{F}}={\mathcal{A}}$, resp. ${\mathcal{C}}={\mathcal{A}}$. We note that any projective or injective abelian model structure on ${\mathcal{A}}$ is hereditary, c.f. [7, Corollary 1.1.12].
@@
The arguments given in [15, Theorem 9.3] for Gorenstein projective/injective model structures carry over mutatis mutandis to this more general situation. ∎
Since the Gorenstein projective, resp. Gorenstein injective, model structure on ${\operatorname{Mod}}(R)$ is projective, resp. injective, it is hereditary. In particular, it is stable which is also proved directly in [15, Theorem 9.3]. Hence we see that the homotopy category ${\operatorname{Ho}}(R)$ has the structure of a triangulated category as described earlier. We note that since every object is cofibrant in the Gorenstein injective model structure the triangles (2.8) are obtained for any short exact sequence in ${\operatorname{Mod}}(R)$.
@@ 3.1. Parabolic induction and recollement: one parabolic
The above setup applies to $S={\mathcal{H}}_{M}$ ($M$ some standard Levi subgroup), $R={\mathcal{H}}$ and ${\mathsf{G}}\in\{{\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}},{\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}\}$ along with its two adjoints. The left adjoint is exact in either case because ${\operatorname{Loc}}_{M}$ is a localisation functor and the equivalence 2.2 is also exact. We’ve also seen that parabolic (co)induction is fully faithful in 2.5.
Note that the functor ${\mathsf{G}}$ is automatically exact as it is both a left and a right adjoint. But ${\mathsf{R}}$ is only known to be left exact, being a right adjoint. Part (i) of the Lemma below is a generalisation of [19, Lemma 8.1] to our setup:
@@
for all ${\mathfrak{m}}\in{\operatorname{Mod}}(S)$, all $\mathfrak{n}\in{\operatorname{Mod}}(R)$ and all $i\geq 0$. This immediately implies that if ${\operatorname{pd}}_{S}({\mathfrak{m}})<\infty$ then ${\operatorname{pd}}_{R}({\mathsf{G}}({\mathfrak{m}}))<\infty$ by considering their injective dimensions, and it also immediately implies (ii). For the converse of (i), given any ${\mathfrak{m}}^{\prime}\in{\operatorname{Mod}}(S)$, 2.4 and (3.1) together give isomorphisms
@@
This now gives that ${\operatorname{pd}}_{R}({\mathsf{G}}({\mathfrak{m}}))<\infty$ implies ${\operatorname{pd}}_{S}({\mathfrak{m}})<\infty$.
@@
We next show that the Gorenstein projective/injective model structures on ${\operatorname{Mod}}(R)$ are in fact determined by those on ${\operatorname{Mod}}(S)$. Note that since the notion of weak equivalences doesn’t depend on which model structure we choose to work with (c.f. 2.12), there is no ambiguity in the statement of part (iii) of the Proposition below.
@@
With $S$, $R$, ${\mathsf{G}}$ and ${\mathsf{L}}$ as in our setup, we have:
@@
We first show that ${\mathsf{G}}$ both preserves and reflects weak equivalences. Since ${\mathsf{G}}$ is exact, it follows from 2.8(i)-(ii) that we only need to show that ${\mathfrak{m}}\in{\operatorname{Mod}}(S)$ is a trivial object, i.e. has finite projective dimension, if and only if ${\mathsf{G}}({\mathfrak{m}})$ is a trivial object in ${\operatorname{Mod}}(R)$. This holds by 3.3(i) and thus proves our claim. To show (i), we are therefore left to prove that for a morphism $f:{\mathfrak{m}}\to{\mathfrak{m}}^{\prime}$ in ${\operatorname{Mod}}(S)$, $f$ is a fibration if and only if ${\mathsf{G}}(f)$ is a fibration, in the Gorenstein projective model structures on ${\operatorname{Mod}}(S)$ and ${\operatorname{Mod}}(R)$ respectively. But in this case the fibrations are the epimorphisms and the result follows because ${\mathsf{G}}$ is exact and faithful.
To show (ii), we instead need to show that $f$ as above is a fibration, resp. cofibration, if and only if ${\mathsf{G}}(f)$ is a fibration, resp. cofibration, in the Gorenstein injective model structures on ${\operatorname{Mod}}(S)$ and ${\operatorname{Mod}}(R)$ respectively. The case of cofibrations is dual to the above: they are the monomorphisms, so that the claim holds by exactness and faithfulness of ${\mathsf{G}}$. The fibrations on the other hand are precisely the epimorphisms with Gorenstein injective kernels. Now the claim holds instead by exactness of ${\mathsf{G}}$ and 3.3(iii).
@@
Continuing with our setup, we have:
@@
    the two total right derived functors of ${\mathsf{G}}$ (with respect to the Gorenstein projective and Gorenstein injective model structures respectively) and the total left derived functor of ${\mathsf{G}}$ (with respect to the Gorenstein injective model structure) are all naturally isomorphic to ${\operatorname{Ho}}({\mathsf{G}})$. Similarly, the two total left derived functor of ${\mathsf{L}}$ are naturally isomorphic to ${\operatorname{Ho}}({\mathsf{L}})$;
@@
Part (i) is immediate from 2.8. For the rest of this proof, we work with the Gorenstein injective model structure. This suffices to show (iii) because the derived functors of ${\mathsf{G}}$ (resp. ${\mathsf{L}}$) for both model structures can be identified with ${\operatorname{Ho}}({\mathsf{G}})$ (resp. ${\operatorname{Ho}}({\mathsf{L}})$), and moreover an adjoint pair is a Quillen equivalence precisely when the derived functors form an adjoint equivalence of categories (c.f. [14, Proposition 1.3.13]).
(ii) By 2.4 we need to show that the counit of the adjunction $({\operatorname{Ho}}({\mathsf{L}}),{\operatorname{Ho}}({\mathsf{G}}))$ is an isomorphism. By the proof of [14, Proposition 1.3.13], this holds if and only if, for any ${\mathfrak{m}}\in{\operatorname{Mod}}(S)$, the composite
@@
is a weak equivalence. Here $Q_{c}{\mathsf{G}}{\mathfrak{m}}$ denotes a cofibrant replacement of ${\mathsf{G}}{\mathfrak{m}}$ in ${\operatorname{Mod}}(R)$, which comes equipped with a trivial fibration $q_{c}:Q_{c}{\mathsf{G}}{\mathfrak{m}}\to{\mathsf{G}}{\mathfrak{m}}$. But by 3.4(iii) we have that ${\mathsf{L}}q_{c}$ is a weak equivalence and $\varepsilon_{{\mathfrak{m}}}$ is an isomorphism (c.f. 2.4 again), so that our claim holds.
(iii) By the above $\varepsilon_{{\mathfrak{m}}}{\mathsf{L}}q_{c}$ is always a weak equivalence. It therefore follows from [14, Corollary 1.3.16] that $({\mathsf{L}},{\mathsf{G}})$ is a Quillen equivalence if and only if ${\mathsf{L}}$ reflects weak equivalences between cofibrant objects. But in the Gorenstein injective model structure all objects are cofibrant, so the latter condition says that ${\mathsf{L}}$ reflects all weak equivalences. The result now follows from 2.8(ii). ∎
Note that from 3.4(iii) and exactness we get that ${\mathsf{L}}$ preserves both fibrations and trivial fibrations in the Gorenstein projective model structure. This is somewhat unusual for a left Quillen functor. Also, we will see later that the Quillen equivalence as in 3.5(iii) essentially never holds in the case of parabolic (co)induction from a standard Levi subgroup $M$ unless $M=G$, c.f. 3.20. For now we only illustrate that this is clear when $M$ is the maximal split torus.
@@
If $M=T$ is the maximal split torus, then ${\mathcal{H}}_{T}=k[\widetilde{\Lambda}]$ is the group algebra of $\widetilde{\Lambda}=T/(T\cap I)\cong{\mathbb{Z}}^{r_{ss}+r_{Z}}\times T(\mathbb{F}_{q})$. Hence ${\mathcal{H}}_{T}$ is free of finite rank, coprime with $p$, over the Laurent polynomial ring $k[x_{1}^{\pm 1},\ldots,x_{r_{ss}+r_{Z}}^{\pm 1}]$. The latter ring has finite global dimension, and therefore so does ${\mathcal{H}}_{T}$ (c.f. [21, Theorem 7.5.6(iii)]). In particular, the homotopy category ${\operatorname{Ho}}({\operatorname{Mod}}({\mathcal{H}}_{T}))$ is zero. On the other hand, ${\mathcal{H}}$ typically has infinite global dimension. Indeed, when $G$ does not have root system of type $A_{1}\times\cdots\times A_{1}$ or when $r_{ss}>0$ and $q>2$, then ${\mathcal{H}}$ has infinite global dimension and in fact has a simple module ${\mathfrak{m}}$ of infinite projective dimension (c.f. [25, 7.1-7.3]). So, under these hypotheses, ${\operatorname{Ho}}({\mathcal{H}})\neq 0$ and ${\operatorname{Ind}}_{{\mathcal{H}}_{T}}^{{\mathcal{H}}}$ is not a Quillen equivalence. And indeed, for this module ${\mathfrak{m}}$ one automatically has ${\operatorname{pd}}_{{\mathcal{H}}_{T}}({\operatorname{Loc}}_{M}(\mathfrak{n}))<\infty$ even though ${\operatorname{pd}}_{{\mathcal{H}}}(\mathfrak{n})=\infty$.
@@
One direction of (iii) is immediate from the spectral sequence (3.3). So assume now that ${\operatorname{Ext}}_{R}^{j}({\mathsf{G}}(\mathfrak{n}),{\mathfrak{m}})\cong{\operatorname{Ext}}^{j}_{S}(\mathfrak{n},{\mathsf{R}}({\mathfrak{m}}))$ for all $\mathfrak{n}\in{\operatorname{Mod}}(S)$ and all $j\geq 0$. Then combining this isomorphism with (3.4) we obtain
@@
When ${\mathsf{G}}={\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$ or ${\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$ for some proper Levi subgroup $M\neq G$, the essential image of ${\mathsf{G}}$ is not closed under subobjects. Indeed, first let $V={\operatorname{Ind}}_{P}^{G}(\mathbf{1}_{M})\in{\operatorname{Rep}}_{k}^{\infty}(G)$. By standard properties of induction for smooth representations, the trivial representation $\mathbf{1}_{G}$ embeds as a subrepresentation of $V$. By left exactness of the functor $U$ and applying [27, Theorem 1], we see that the trivial character $\mathbf{1}_{{\mathcal{H}}}$ of ${\mathcal{H}}$ embeds as a submodule of ${\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}(\mathbf{1}_{{\mathcal{H}}_{M}})$. However, there does not exist ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}}_{M})$ such that ${\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}({\mathfrak{m}})\cong\mathbf{1}_{{\mathcal{H}}}$. This is because, by [29, Lemmas 3.5 & 3.6], we have an isomorphism ${\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}({\mathfrak{m}})\cong{\mathfrak{m}}^{\oplus n}$ of $k$-vector spaces, where $n=\left|W_{0}\right|/\left|W_{0,M}\right|>1$. So in particular $\dim_{k}{\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}({\mathfrak{m}})>1$ whenever ${\mathfrak{m}}\neq 0$.
@@
Let ${\mathcal{T}}$ be a triangulated category with shift functor $\Sigma:{\mathcal{T}}\to{\mathcal{T}}$. Recall that a subcategory ${\mathcal{S}}$ of ${\mathcal{T}}$ is called a triangulated subcategory if $\Sigma$ restricts to an autoequivalence of ${\mathcal{S}}$, if ${\mathcal{S}}$ is strictly full (meaning any object of ${\mathcal{T}}$ isomorphic to an object of ${\mathcal{S}}$ lies in ${\mathcal{S}}$) and if whenever $X\to Y\to Z\to\Sigma X$ is a distinguished triangle in ${\mathcal{T}}$ with two of $X,Y,Z$ lying in ${\mathcal{S}}$, then so does the third one. A triangulated subcategory ${\mathcal{S}}$ is called thick if, furthermore, it is closed under direct summands. If ${\mathcal{T}}^{\prime}$ is another triangulated category with shift functor $\Sigma^{\prime}$, an additive functor ${\mathsf{F}}:{\mathcal{T}}\to{\mathcal{T}}^{\prime}$ is called triangulated if ${\mathsf{F}}\Sigma\cong\Sigma^{\prime}{\mathsf{F}}$ and ${\mathsf{F}}$ sends distinguished triangles to distinguished triangles. Note that the essential image of a full triangulated functor is a triangulated subcategory. Moreover, if ${\mathcal{S}}$ is a triangulated subcategory then it inherits from ${\mathcal{T}}$ a triangulated structure and the inclusion ${\mathcal{S}}\to{\mathcal{T}}$ is a triangulated functor.
Suppose now that ${\mathcal{T}}$ contains all small coproducts. For example, the homotopy category of any Gorenstein ring has all small coproducts since Gorenstein projective modules are closed under arbitrary direct sums. Then recall that a triangulated subcategory ${\mathcal{S}}$ is called localising if it is closed under arbitrary small coproducts. All such subcategories are automatically thick, c.f. [22, Proposition 1.6.8].
In general there is a Verdier quotient ${\mathcal{T}}/{\mathcal{S}}$, for ${\mathcal{T}}$ a triangulated category and ${\mathcal{S}}$ a triangulated subcategory, which is itself a triangulated category equipped with a triangulated quotient functor $\pi:{\mathcal{T}}\to{\mathcal{T}}/{\mathcal{S}}$, and this construction satisfies a suitable universal property (c.f. [28, §II.2.1-II.2.3]).
@@
We briefly describe some properties implied by this definition (c.f. [20, §4.9-4.13]). First, ${\mathcal{T}}^{\prime}$ identifies with a thick subcategory of ${\mathcal{T}}$ via ${\mathsf{G}}$ such that ${\mathcal{T}}^{\prime\prime}\simeq{\mathcal{T}}/{\mathcal{T}}^{\prime}$. Moreover, the adjunction morphisms give for any $X\in{\mathcal{T}}$ two functorial distinguished triangles
$\mathsf{G}\mathsf{R}(X)\to X\to\mathsf{P}\mathsf{I}(X)\to\Sigma(\mathsf{G}\mathsf{R}(X))$

### v4

@@ 2.4. Recollections on abelian model categories
By combining 2.14 and 2.15, we see that for ${\mathfrak{m}},\mathfrak{n}\in{\operatorname{Mod}}(R)$ with ${\mathfrak{m}}$ Gorenstein projective and for $\theta:\mathfrak{p}\to\mathfrak{n}$ as above, $[{\mathfrak{m}},\mathfrak{n}]_{R}\cong{\operatorname{coker}}(\theta\circ(-))\cong{\operatorname{Ext}}^{1}_{S}({\mathfrak{m}},\ker(\theta))$. This generalises as follows. We say that an abelian model structure $({\mathcal{C}},{\mathcal{F}},{\mathcal{W}})$ on an abelian category ${\mathcal{A}}$ is projective, resp. injective, if ${\mathcal{F}}={\mathcal{A}}$, resp. ${\mathcal{C}}={\mathcal{A}}$. We note that any projective or injective abelian model structure on ${\mathcal{A}}$ is hereditary, cf. [7, Corollary 1.1.12].
@@
The arguments given in [17, Theorem 9.3] for Gorenstein projective/injective model structures carry over mutatis mutandis to this more general situation. ∎
Since the Gorenstein projective, resp. Gorenstein injective, model structure on ${\operatorname{Mod}}(R)$ is projective, resp. injective, it is hereditary. In particular, it is stable which is also proved directly in [17, Theorem 9.3]. Hence we see that the homotopy category ${\operatorname{Ho}}(R)$ has the structure of a triangulated category as described earlier. We note that since every object is cofibrant in the Gorenstein injective model structure the triangles (2.11) are obtained for any short exact sequence in ${\operatorname{Mod}}(R)$.
@@ 3.1. Parabolic induction and recollement: one parabolic
The above setup applies to $S={\mathcal{H}}_{M}$ ($M$ some standard Levi subgroup), $R={\mathcal{H}}$ and ${\mathsf{G}}\in\{{\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}},{\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}\}$ along with its two adjoints. The left adjoint is exact in either case because ${\operatorname{Loc}}_{M}$ is a localisation functor and the equivalence in 2.2 is also exact. We’ve also seen that parabolic (co)induction is fully faithful in 2.5.
Note that the functor ${\mathsf{G}}$ is automatically exact as it is both a left and a right adjoint. But ${\mathsf{R}}$ is only known to be left exact, being a right adjoint. Part (i) of the Lemma below is a generalisation of [21, Lemma 8.1] to our setup:
@@
for all ${\mathfrak{m}}\in{\operatorname{Mod}}(S)$, all $\mathfrak{n}\in{\operatorname{Mod}}(R)$ and all $i\geq 0$. This gives that ${\operatorname{id}}_{R}({\mathsf{G}}({\mathfrak{m}}))\leq{\operatorname{id}}_{S}({\mathfrak{m}})$ and also immediately implies (ii). Furthermore, given any ${\mathfrak{m}}^{\prime}\in{\operatorname{Mod}}(S)$, 2.4 and (3.1) together give isomorphisms
@@
We deduce immediately that ${\operatorname{pd}}_{S}({\mathfrak{m}})\leq{\operatorname{pd}}_{R}({\mathsf{G}}({\mathfrak{m}}))$. Since $R$ and $S$ are assumed to be Gorenstein, finite injective dimension coincides with finite projective dimension for modules over both $R$ and $S$, and (i) follows from the above two inequalities.
@@
We next show that the Gorenstein projective/injective model structures on ${\operatorname{Mod}}(S)$ are in fact determined by those on ${\operatorname{Mod}}(R)$. Note that since the notion of weak equivalences doesn’t depend on which model structure we choose to work with (cf. 2.12), there is no ambiguity in the statement of part (iii) of the Proposition below.
@@
With $S$, $R$, ${\mathsf{G}}$ and ${\mathsf{L}}$ as in 3.1, we have:
@@
We fix a morphism $f$ in ${\operatorname{Mod}}(S)$. Our first goal is to establish the following:
  • (A)
    $f$ is a fibration if and only if ${\mathsf{G}}(f)$ is a fibration, in the Gorenstein projective model structures on ${\operatorname{Mod}}(S)$ and ${\operatorname{Mod}}(R)$; and
  • (B)
    $f$ is a fibration, resp. cofibration, if and only if ${\mathsf{G}}(f)$ is a fibration, resp. cofibration, in the Gorenstein injective model structures on ${\operatorname{Mod}}(S)$ and ${\operatorname{Mod}}(R)$.
To get (A), we simply observe that fibrations are epimorphisms and the statement holds because ${\mathsf{G}}$ is exact and faithful. To get (B), the case of cofibrations is proved completely dually to (A): the cofibrations are the monomorphisms, so that the claim holds by exactness and faithfulness of ${\mathsf{G}}$. The fibrations on the other hand are precisely the epimorphisms with Gorenstein injective kernels. Now the claim for that case holds by exactness and faithfulness of ${\mathsf{G}}$ and by 3.3(iii).
To prove (i)-(ii), it now suffices to show that $f$ is a weak equivalence if and only if ${\mathsf{G}}(f)$ is a weak equivalence. First note that ${\mathsf{G}}$ preserves trivial objects, i.e.  modules of finite projective dimension, by 3.3(i). Since ${\mathsf{G}}$ is exact, it now follows from 2.8(i) that ${\mathsf{G}}$ preserves weak equivalences. In particular, by (A)-(B) above we have that ${\mathsf{G}}$ is right Quillen for both the Gorenstein projective and injective model structures, and left Quillen for the Gorenstein injective model structure. Hence, we may apply 2.8(ii) which, together with 3.3(i) again, gives us that ${\mathsf{G}}$ reflects weak equivalences as required.
@@
Continuing with $S$, $R$, ${\mathsf{G}}$ and ${\mathsf{L}}$ as in 3.1, we have:
@@
    the two total right derived functors of ${\mathsf{G}}$ (with respect to the Gorenstein projective and Gorenstein injective model structures respectively) and the total left derived functor of ${\mathsf{G}}$ (with respect to the Gorenstein injective model structure) are all naturally isomorphic to ${\operatorname{Ho}}({\mathsf{G}})$. Similarly, the two total left derived functors of ${\mathsf{L}}$ are naturally isomorphic to ${\operatorname{Ho}}({\mathsf{L}})$;
@@
Part (i) is immediate from 2.8. For the rest of this proof, we work with the Gorenstein injective model structure. This suffices to show (iii) because the derived functors of ${\mathsf{G}}$ (resp. ${\mathsf{L}}$) for both model structures can be identified with ${\operatorname{Ho}}({\mathsf{G}})$ (resp. ${\operatorname{Ho}}({\mathsf{L}})$), and moreover an adjoint pair is a Quillen equivalence precisely when the derived functors form an adjoint equivalence of categories (cf. [16, Proposition 1.3.13]).
(ii) By 2.4 we need to show that the counit of the adjunction $({\operatorname{Ho}}({\mathsf{L}}),{\operatorname{Ho}}({\mathsf{G}}))$ is an isomorphism. By the proof of [16, Proposition 1.3.13], this holds if and only if, for any ${\mathfrak{m}}\in{\operatorname{Mod}}(S)$, the composite
@@
is a weak equivalence. Here $Q_{c}{\mathsf{G}}{\mathfrak{m}}$ denotes a cofibrant replacement of ${\mathsf{G}}{\mathfrak{m}}$ in ${\operatorname{Mod}}(R)$, which comes equipped with a trivial fibration $q_{c}:Q_{c}{\mathsf{G}}{\mathfrak{m}}\to{\mathsf{G}}{\mathfrak{m}}$. But by 3.4(iii) we have that ${\mathsf{L}}q_{c}$ is a weak equivalence and $\varepsilon_{{\mathfrak{m}}}$ is an isomorphism (cf. 2.4 again), so that our claim holds.
(iii) By the above $\varepsilon_{{\mathfrak{m}}}{\mathsf{L}}q_{c}$ is always a weak equivalence. It therefore follows from [16, Corollary 1.3.16] that $({\mathsf{L}},{\mathsf{G}})$ is a Quillen equivalence if and only if ${\mathsf{L}}$ reflects weak equivalences between cofibrant objects. But in the Gorenstein injective model structure all objects are cofibrant, so the latter condition says that ${\mathsf{L}}$ reflects all weak equivalences. The result now follows from 2.8(ii). ∎
Note that from 3.4(iii) and exactness we get that ${\mathsf{L}}$ preserves both fibrations and trivial fibrations in the Gorenstein projective model structure. This is somewhat unusual for a left Quillen functor. Also, we will see later that the Quillen equivalence as in 3.5(iii) essentially never holds in the case of parabolic (co)induction from a standard Levi subgroup $M$ unless $M=G$, cf. 3.20. For now we only illustrate that this is clear when $M$ is the maximal split torus.
@@
If $M=T$ is the maximal split torus, then ${\mathcal{H}}_{T}=k[\widetilde{\Lambda}]$ is the group algebra of $\widetilde{\Lambda}=T/(T\cap I)\cong{\mathbb{Z}}^{r_{ss}+r_{Z}}\times T(\mathbb{F}_{q})$. Hence ${\mathcal{H}}_{T}$ is free of finite rank, coprime with $p$, over the Laurent polynomial ring $k[x_{1}^{\pm 1},\ldots,x_{r_{ss}+r_{Z}}^{\pm 1}]$. The latter ring has finite global dimension, and therefore so does ${\mathcal{H}}_{T}$ (cf. [23, Theorem 7.5.6(iii)]). In particular, the homotopy category ${\operatorname{Ho}}({\operatorname{Mod}}({\mathcal{H}}_{T}))$ is zero. On the other hand, ${\mathcal{H}}$ typically has infinite global dimension. Indeed, when $G$ does not have root system of type $A_{1}\times\cdots\times A_{1}$ or when $r_{ss}>0$ and $q>2$, then ${\mathcal{H}}$ has infinite global dimension and in fact has a simple module ${\mathfrak{m}}$ of infinite projective dimension (cf. [27, 7.1-7.3]). So, under these hypotheses, ${\operatorname{Ho}}({\mathcal{H}})\neq 0$ and ${\operatorname{Ind}}_{{\mathcal{H}}_{T}}^{{\mathcal{H}}}$ is not a Quillen equivalence. And indeed, for this module ${\mathfrak{m}}$ one automatically has ${\operatorname{pd}}_{{\mathcal{H}}_{T}}({\operatorname{Loc}}_{M}({\mathfrak{m}}))<\infty$ even though ${\operatorname{pd}}_{{\mathcal{H}}}({\mathfrak{m}})=\infty$.
@@
For (iii), the spectral sequence (3.3) immediately implies that the isomorphism of Ext groups holds when ${\mathfrak{m}}$ is is ${\mathsf{R}}$-acyclic. Conversely, assume that ${\operatorname{Ext}}_{R}^{j}({\mathsf{G}}(\mathfrak{n}),{\mathfrak{m}})\cong{\operatorname{Ext}}^{j}_{S}(\mathfrak{n},{\mathsf{R}}({\mathfrak{m}}))$ for all $\mathfrak{n}\in{\operatorname{Mod}}(S)$ and all $j\geq 0$. Then combining this isomorphism with (3.4) we obtain
@@
When ${\mathsf{G}}={\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$ or ${\operatorname{Coind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}$ for some proper Levi subgroup $M\neq G$, the essential image of ${\mathsf{G}}$ is not closed under subobjects. Indeed, first let $V={\operatorname{Ind}}_{P}^{G}(\mathbf{1}_{M})\in{\operatorname{Rep}}_{k}^{\infty}(G)$. By standard properties of induction for smooth representations, the trivial representation $\mathbf{1}_{G}$ embeds as a subrepresentation of $V$. By left exactness of the functor $U$ and applying [30, Theorem 1], we see that the trivial character $\mathbf{1}_{{\mathcal{H}}}$ of ${\mathcal{H}}$ embeds as a submodule of ${\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}(\mathbf{1}_{{\mathcal{H}}_{M}})$. However, there does not exist ${\mathfrak{m}}\in{\operatorname{Mod}}({\mathcal{H}}_{M})$ such that ${\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}({\mathfrak{m}})\cong\mathbf{1}_{{\mathcal{H}}}$. This is because, by [32, Lemmas 3.5 & 3.6], we have an isomorphism ${\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}({\mathfrak{m}})\cong{\mathfrak{m}}^{\oplus n}$ of $k$-vector spaces, where $n=\left|W_{0}\right|/\left|W_{0,M}\right|>1$. So in particular $\dim_{k}{\operatorname{Ind}}_{{\mathcal{H}}_{M}}^{{\mathcal{H}}}({\mathfrak{m}})>1$ whenever ${\mathfrak{m}}\neq 0$.
@@
Let ${\mathcal{T}}$ be a triangulated category with shift functor $\Sigma:{\mathcal{T}}\to{\mathcal{T}}$. Recall that a subcategory ${\mathcal{S}}$ of ${\mathcal{T}}$ is called a triangulated subcategory if $\Sigma$ restricts to an autoequivalence of ${\mathcal{S}}$, if ${\mathcal{S}}$ is strictly full (meaning any object of ${\mathcal{T}}$ isomorphic to an object of ${\mathcal{S}}$ lies in ${\mathcal{S}}$) and if whenever $X\to Y\to Z\to\Sigma X$ is a distinguished triangle in ${\mathcal{T}}$ with two of $X,Y,Z$ lying in ${\mathcal{S}}$, then so does the third one. A subcategory ${\mathcal{S}}$ is called thick if it is triangulated and closed under direct summands. If ${\mathcal{T}}^{\prime}$ is another triangulated category with shift functor $\Sigma^{\prime}$, an additive functor ${\mathsf{F}}:{\mathcal{T}}\to{\mathcal{T}}^{\prime}$ is called triangulated if ${\mathsf{F}}\Sigma\cong\Sigma^{\prime}{\mathsf{F}}$ and ${\mathsf{F}}$ sends distinguished triangles to distinguished triangles. Note that the essential image of a full triangulated functor is a triangulated subcategory. Moreover, if ${\mathcal{S}}$ is a triangulated subcategory then it inherits from ${\mathcal{T}}$ a triangulated structure and the inclusion ${\mathcal{S}}\to{\mathcal{T}}$ is a triangulated functor.
Suppose now that ${\mathcal{T}}$ contains all small coproducts. For example, the homotopy category of any Gorenstein ring has all small coproducts since Gorenstein projective modules are closed under arbitrary direct sums. Then recall that a triangulated subcategory ${\mathcal{S}}$ is called localising if it is closed under arbitrary small coproducts. All such subcategories are automatically thick, cf. [24, Proposition 1.6.8].
In general there is a Verdier quotient ${\mathcal{T}}/{\mathcal{S}}$, for ${\mathcal{T}}$ a triangulated category and ${\mathcal{S}}$ a triangulated subcategory, which is itself a triangulated category equipped with a triangulated quotient functor $\pi:{\mathcal{T}}\to{\mathcal{T}}/{\mathcal{S}}$, and this construction satisfies a suitable universal property (cf. [31, §II.2.1-II.2.3]).
@@
We briefly describe some properties implied by this definition (cf. [22, §4.9-4.13]). First, ${\mathcal{T}}^{\prime}$ identifies with a thick subcategory of ${\mathcal{T}}$ via ${\mathsf{G}}$ such that ${\mathcal{T}}^{\prime\prime}\simeq{\mathcal{T}}/{\mathcal{T}}^{\prime}$. Moreover, the adjunction morphisms give for any $X\in{\mathcal{T}}$ two functorial distinguished triangles
(3.5) $\mathsf{G}\mathsf{R}(X)\to X\to\mathsf{P}\mathsf{I}(X)\to\Sigma(\mathsf{G}\mathsf{R}(X))$